# Time-resolved diamond magnetic microscopy of superparamagnetic iron-oxide nanoparticles

### v1

@@ Appendix III.2 SEM imaging
After SPIONs are deposited on the diamond, a region near the center of the diamond surface is selected for SEM imaging. SPIONs are imaged with an FEI Helios NanoLab 650 scanning electron microscope (SEM) to locate single, isolated SPIONs. Images are taken with a magnification of 10000×10000$\times$10000 ×, and each image region of interest (ROI) is approximately $42\times 28~{}{\rm\upmu m^{2}}$.
@@
For the field of view (FOV) presented in the main text, six of these SEM images were acquired, with some overlap between adjacent images, and stitched together using the software FIJI [76], forming a $2\times 3$ grid of images. The ROI of the stitched image was approximately $75\times 75~{}\upmu$m2. (SEM images were taken over a second ROI of similar size elsewhere on the diamond surface, and a similar SPION density was observed.) The stitched image is then overlaid onto the magnetic microscope image, rotated and resized to match the magnetic image field of view. A small degree of distortion of the SEM image is used to precisely align magnetic dipole patterns with corresponding SPION SEM features. During all image transformations, a fiducial marker milled into the diamond surface (visible in both the SEM and magnetic images) is used to ensure that the overlay is accurate to within $\lesssim 1~{}{\rm\upmu m}$. The final SEM image is then cropped to match the magnetic image FOV.
@@
Following the initial processing, we used Wolfram Mathematica’s morphological image analysis function ComponentMeasurements to locate the pixel coordinates of the centroid of each single SPION or SPION aggregate. A list of these image coordinates was generated, together with a corresponding list identifying each feature as a single SPION, SPION aggregate, or unidentified non-SPION particulate. Additionally, it was noted whether each feature was ${\gtrsim}1.5~{}{\rm\upmu m}$ or ${\gtrsim}1~{}{\rm\upmu m}$ from its nearest neighboring feature. (Distances within the SEM image were quantified by measuring the number of pixels corresponding to the scale bar generated by the SEM. Some error is introduced by the rotation and distortions described above, as well as sample charging during image acquisition, but the error is small enough that feature co-localization with magnetic images was still possible.) Figure A8 shows a magnetic image taken with an applied field of $B_{0}=37.2~{}{\rm mT}$. The image is annotated with boxes around the magnetic features corresponding to single SPIONs that are each $\gtrsim 1~{}{\rm\upmu m}$ from their nearest neighbor as determined by SEM imaging.
@@
We imaged more than three ${\sim}75\times 75~{}{\rm\upmu m^{2}}$ ROIs with the SEM, taken over several repetitions of the same SPION drop-casting process. Some of these regions were not studied with the magnetic microscope due to a low number of isolated single SPIONs (owing to a SPION-suspension number density that was either too high or too low). For other regions, magnetic images were taken over a range of applied fields, but the studies described in the main text were unable to be completed due to degradation of the optical adhesive under the diamond after exposure to the $532~{}{\rm nm}$ laser light for ${\gtrsim}4$ weeks.
@@ Appendix IV.1 Concept
where $C$ is the contrast, $\Gamma$ is the full-width at half-maximum (FWHM) linewidth, $F_{0}$ is the off-resonance fluorescence intensity, and $\Delta_{c}$ is the resonance’s central frequency with respect to the scan’s center. For a single pixel, the typical count rate was $F_{0}\approx 40~{}{\rm kilocounts}$ for a $5.2~{}{\rm ms}$ exposure time, the ODMR contrast was $C\approx 0.03$, and the FWHM linewidth was $\Gamma\approx 12~{}{\rm MHz}$.
Appendix IV.2 Parallel ODMR fitting
@@ Appendix IV.3 Sequences and timing
The summand in Eq. (AVI-4) is a convolution of the $x$-component of the SPION dipole field, $B_{\rm dip,x}(x,y)$, with the microscope optical point-spread function, ${\rm PSF}(x,y)$. This expression is averaged over 5 different $z$ depths, ($z_{i}=-\{60,90,120,150,180\}~{}{\rm nm}$) selected to represent the relevant NV layer distribution, as described in Appendix II. The $60~{}{\rm nm}$ standoff is chosen to account for the SPION radius (${\sim}15~{}{\rm nm}$), a thin hydrocarbon layer between SPION and diamond(${\sim}10~{}{\rm nm}$), and a (${\sim}15~{}{\rm nm}$) depleted layer near the diamond surface where implanted nitrogen is not efficiently converted to negatively-charged NV centers [36]. An additional consideration in choosing the closest representative NV layer is the expectation that NV centers closer than ${\sim}60~{}{\rm nm}$ to a SPION center exhibit large enough ODMR line shifts that their contribution to the voxel-averaged ODMR lineshape is inadequately captured by a least-squares Lorentzian fit.
@@
Figure A11: Easy axis angle. An approximate value of $\cos{\theta}$ ) is given by $m_{x}/|\vec{m}|$, where $\theta$ is the angle of the easy axis with respect to the applied field direction and $\vec{m}$ is extract from fits to SPIONs at the lowest practical applied field ($B_{0}=1.5~{}{\rm mT}$). A scatter plot of $m_{x}/|\vec{m}|$ vs. $B_{c}$ is shown for ${\sim}95$ SPIONs. As the standard expression for characteristic field is $B_{c}=k_{B}T/(\mu\cos{\theta})$ ), the approximately inverse relationship observed in this plot is expected.
@@
Appendix VIII.1 Sample preparation and measurement
A Quantum Design MPMS-3 SQUID magnetometer was used to characterize SPIONs of the same batch (SOR30, Ocean Nanotech) as those imaged by diamond magnetic microscopy. Samples were prepared by suspending SPIONs in hexane, sonicating for approximately $10$ minutes to discourage particle aggregation, and depositing the suspension on a cotton swab. The cotton swab was fixed in a plastic straw sample holder for measurement inside the MPMS. An alternative sample preparation we used for bulk SPION characterization involved freeze-drying a SPION suspension in mannitol [78].
@@ Appendix VIII.2 Field-dependent magnetization curves
The mean of the single-SPION $m_{\rm sat}$ distribution [$\overline{m}_{\rm sat}=0.44~{}{\rm A{\cdot}nm^{2}}$, see Fig. 3(c)] can also be compared with literature values. Wu et al. [52] reported a saturation magnetic moment of $\overline{m}_{\rm sat}=2.5~{}{\rm A{\cdot}nm^{2}}$ per nanoparticle for 30-nm-diameter Ocean Nanotech SHA-30 SPIONs, and $\overline{m}_{\rm sat}=1.6~{}{\rm A{\cdot}nm^{2}}$ per nanoparticle for 25-nm-diameter Ocean Nanotech SHA-25 SPIONs. Ahrentorp et al. [51] reported $\overline{m}_{\rm sat}=2\mbox{-}3.4~{}{\rm A{\cdot}nm^{2}}$ for 25-nm SPIONs suspended in liquid. Mosavian et al. [36] used a super-resolution diamond magnetic microscopy technique to record $m_{x}=0.86~{}{\rm A{\cdot}nm^{2}}$ for a single Ocean Nanotech SOR-30 SPION at an applied field of $30~{}{\rm mT}\,\hat{x}$. Our diamond magnetic microscopy measurements on 30-nm SPIONs yield $\overline{m}_{\rm sat}=0.44~{}{\rm A{\cdot}nm^{2}}$, which is somewhat smaller than these values.
There are several possible reasons why our method might record spuriously lower measured magnetic moments. Diamond magnetic microscopy can be affected by pixel cross-talk in which signals from distant NV centers contribute to the ODMR spectrum of a camera pixel near a SPION [36, 81]. Additionally, it is possible that the distribution of NV depths in the diamond is broader and deeper than estimated by SRIM; for example, due to ion channeling effects. This could be prevented by using delta-doped NV layers [66, 67, 68]. Furthermore, our estimation of $\vec{m}$ assumes a 30 nm magnetic core, but this was not directly measured.
@@ Appendix VIII.3 AC susceptibility
The bulk AC susceptibility data is approximately log-normally distributed. The peak of the function fit to the 400 K data in A13(a) is $f_{\rm max}\approx 80~{}{\rm Hz}$, which corresponds to a modal value of $\tau_{N}\approx 2~{}{\rm ms}$. Taking the holding-field dependence $\tau_{N}=\tau_{N,0}\exp(-B_{\rm hold}/B_{N})$, as observed in Fig. 5(b) with $B_{N}\approx 0.7~{}{\rm mT}$, this translates to a modal value $\tau_{N}\approx 0.1~{}{\rm ms}$ at $B_{\rm hold}=2~{}{\rm mT}$. This peak value of $\tau_{N}$, together with a $\sim 6$-decade FWHM, give a log-normal distribution that approximately matches the experimental histogram data in Fig. 5(b) ($N=83$ for $\tau_{N}<0.06~{}{\rm s}$, $N=11$ for $0.06~{}{\rm s}<\tau_{N}<10~{}{\rm s}$, and $N=3$ for $\tau_{N}>10~{}{\rm s}$).
@@
Figures A15-A19 shows plots of the fitted values for $m_{x}$, $m_{y}$, and $m_{z}$ versus $B_{0}$ for each of the 101 SPIONs studied in Fig. 3 of the main text. The $m_{x}(B_{0})$ curves are fit to a Langevin function, see Eq. (1) of the main text. There are five features (numbered 66, 109, 117, 140, and 220) that are designated as outliers due to having fit results $B_{c}>7~{}{\rm mT}$ and $m_{\rm sat}>1.2~{}{\rm A{\cdot}nm^{2}}$. These outliers exhibit little-to-no saturation at the highest applied fields, so their Langevin fits are ambiguous.
@@
Figures A20-A22 show plots of $\Delta m(t)$ for 93 of the SPIONs in the field of view.

### v2

@@ Appendix III.2 SEM imaging
After SPIONs are deposited on the diamond, a region near the center of the diamond surface is selected for SEM imaging. SPIONs are imaged with an FEI Helios NanoLab 650 SEM to locate single, isolated SPIONs. To mitigate charging during SEM imaging, a copper clip mount was used, with the clip pressing on the diamond face itself. The current and accelerating voltage were lowered to $50~{}{\rm pA}$ and $5~{}{\rm kV}$ respectively. Even with these measures, charging was found to be highly variable from diamond to diamond. The diamond used for all of the measurements reported in the manuscript (described in Appendix II) exhibited sufficiently low charging for imaging, allowing for faithful co-localization with diamond magnetic images after correcting for a small degree of distortion. Images were taken with a magnification of 10000×10000$\times$10000 ×, and each image region of interest (ROI) is approximately $42\times 28~{}{\rm\upmu m^{2}}$.
@@
For the field of view (FOV) presented in the main text, six of these SEM images were acquired, with some overlap between adjacent images, and stitched together using the software FIJI [80], forming a $2\times 3$ grid of images. The ROI of the stitched image was approximately $75\times 75~{}{\rm\upmu m^{2}}$. SEM images were taken over a second ROI of similar size elsewhere on the diamond surface, and a similar SPION density was observed. The stitched image is then overlaid onto the magnetic microscope image, rotated and resized to match the magnetic image field of view. A small degree of distortion of the SEM image is used to precisely align magnetic dipole patterns with corresponding SPION SEM features. During all image transformations, a fiducial marker milled into the diamond surface (visible in both the SEM and magnetic images) is used to ensure that the overlay is accurate to within $\lesssim 1~{}{\rm\upmu m}$. The final SEM image is then cropped to match the magnetic image FOV.
@@
Following the initial processing, we used Wolfram Mathematica’s morphological image analysis function, “ComponentMeasurements”, to locate the pixel coordinates of the centroid of each single SPION or SPION aggregate. A list of these image coordinates was generated, together with a corresponding list identifying each feature as a single SPION, SPION aggregate, or unidentified non-SPION particulate. Additionally, it was noted whether each feature was ${\gtrsim}1.5~{}{\rm\upmu m}$ or ${\gtrsim}1~{}{\rm\upmu m}$ from its nearest neighboring feature. (Distances within the SEM image were quantified by measuring the number of pixels corresponding to the scale bar generated by the SEM. Some error is introduced by the rotation and distortions described above, as well as sample charging during image acquisition, but the error is small enough that feature co-localization with magnetic images was still possible.)
Figure A8 shows a magnetic image taken with an applied field of $B_{0}=37.2~{}{\rm mT}$. The image is annotated with boxes around the magnetic features corresponding to single SPIONs that are each $\gtrsim 1~{}{\rm\upmu m}$ from their nearest neighbor as determined by SEM imaging.
@@
We imaged more than three ${\sim}75\times 75~{}{\rm\upmu m^{2}}$ ROIs with the SEM, taken over several repetitions of the same SPION drop-casting process. Some of these regions were not studied with the magnetic microscope due to a low number of isolated single SPIONs (owing to a SPION-suspension number density that was either too high or too low). For other regions, magnetic images were taken over a range of applied fields, but the full set of studies described in the main text were unable to be completed due to degradation of the optical adhesive under the diamond after exposure to the $532~{}{\rm nm}$ laser light for ${\gtrsim}4$ weeks.
Figure A9: Diamond surface before and after drop-casting. (a) SEM image of the diamond surface after cleaning but before drop-casting SPIONs. (b) SEM image of the diamond surface after SPION deposition. The scale bars in both (a) and (b) are $20~{}{\rm\upmu m}$.
In SEM images with SPIONs on diamond, for example Fig. A7, there are pronounced contrast “blemishes” of varying size. To investigate their origin, we acquired SEM images of the clean diamond surface prior to depositing SPIONs on the diamond surface. The diamond was cleaned in a solution of 1:1:1 nitric:sulfuric:perchloric acid at $\sim 200\degree$ C for several hours followed by rinsing with deionized water and then isopropyl alcohol. Figure A9(a) shows an SEM image of the diamond, taken after this cleaning procedure. Next we drop-cast SPIONs on to the diamond surface and acquired a second image of a similar region of the diamond, Fig. A9(b). Comparison of these images suggests that most of the dark blemishes are caused by residues from the hexane in the initial SPION suspension. Some features may also be due to inadvertent contact between the diamond and its container.
Appendix III.3 TEM imaging
Figure A10: Transmission electron microscopy (TEM) images of SPIONs. (a) Low-magnification TEM image of SPIONs used to determine shape and size distribution. Inset: selected area electron diffraction pattern consistent with face-centered cubic Fe3O4. (b) High-resolution transmission electron microscopy image of a SPION showing indexed lattice fringes. The relative orientation of the planes is inconsistent with the SPION being a single crystal.
The standard drop-casting method, as described in [81], was used to prepare the SPIONs for transmission electron microscopy (TEM) analysis. This process involved extracting a small volume of the SPIONs dispersed in chloroform using a micropipette and depositing it onto a TEM copper grid. The sample was then allowed to air dry for 3 hours to ensure the complete evaporation of the solvent before TEM analysis. TEM analysis was conducted using the FEI Tecnai Osiris operating at 200 kV. As seen in Fig. A10(a), the SPIONs exhibit a spherical morphology with a nearly uniform particle diameter. Using an image segmentation program, we determined the mean particle diameter to be $30~{}{\rm nm}$, with a standard deviation of ${\sim}4~{}{\rm nm}$. A selected area electron diffraction pattern, inset of Fig. A10(a), reveals distinct diffraction rings indexing to face-centered cubic Fe3O4. High-resolution images, Fig. A10(b), show lattice spacings consistent with the $\{220\}$ and $\{311\}$ interplanar spacings for face-centered cubic Fe3O4 crystal structure, corroborating the diffraction results. The high-resolution TEM images also reveal evidence that the SPIONS were polycrystalline. Lattice fringes from a given plane do not extend across the diameter of the nanoparticle, and the observed $(220)$ and $(311)$ planes in Fig. A10(b) are nearly parallel, which structurally cannot occur in a single crystal.
@@ Appendix IV.1 Concept
where $C$ is the ODMR contrast, $\Gamma$ is the full-width at half-maximum (FWHM) linewidth, $F_{0}$ is the off-resonance fluorescence intensity, and $\Delta_{c}$ is the resonance’s central frequency with respect to the scan’s center. For a single pixel, the typical count rate was $F_{0}\approx 40~{}{\rm kilocounts}$ for a $5.2~{}{\rm ms}$ exposure time, the ODMR contrast was $C\approx 0.03$, and the FWHM linewidth was $\Gamma\approx 12~{}{\rm MHz}$.
Appendix IV.2 Sensitivity enhancement
In the main text, we report a ${\sim}1.4$-fold improvement in sensitivity when using dual-resonance magnetometry, as compared to the typical single-resonance approach. This improvement is based on the observed change in the ODMR spectrum upon driving both resonances simultaneously. Here, we discuss factors contributing to the observed increase in ODMR contrast.
The minimum detectable magnetic field is directly proportional to the ratio $\Gamma/C$, where $\Gamma$ is the ODMR FWHM linewidth and $C$ is the ODMR contrast [82]. In our experiments, we first minimized this ratio (by adjusting microwave power) under single-resonance driving. Next, we did the same optimization under dual-resonance driving. We observed that the optimized $\Gamma/C$ ratio decreased by a factor of ${\sim}1.4$ when we applied dual resonance (compared to single resonance), indicating a ${\sim}1.4$-fold improvement in sensitivity. Specifically, we observed that $C$ increased by a factor of ${\sim}1.4$ while $\Gamma$ hardly changed. This improvement in sensitivity turned out to be quite helpful in reducing the experimental integration times.
In an ideal Ramsey-type pulsed ODMR measurement, there should be no improvement in fluorescence contrast when using dual-resonance excitation [83]. The situation in continuous-wave ODMR is more complicated, as there is a rich interplay between the optical excitation rate ($I_{0}$), microwave Rabi frequency ($\Omega_{R}$), NV spin dephasing rate ($1/T_{2}^{\ast}$), and NV longitudinal spin relaxation rate ($1/T_{1}$[84, 85]. In the limit $1/T_{2}^{\ast}\gg\Omega_{R}\gg I_{0}\gg 1/T_{1}$, we can apply a classical rate equation model to gain intuition about the dynamics. In this case, as shown in [46], under dual resonance conditions, the population in $m_{s}=0$ approaches $1/3$. On the other hand, in the single resonance case, the population in $m_{s}=0$ would be ${\sim}1/2$. Thus, in this toy model, the contrast improves by a factor of ${\sim}4/3$. Another extreme is the situation where the NV dephasing rate is low ($T_{2}^{\ast}$ is long) and $\Omega_{R}$ is relatively high, as explored in Ref. [48]. There, it was observed that the contrast under dual resonance excitation can actually decrease (compared to single resonance), which was attributed to complicated dynamics that counteract population trapping in coherent dark states.
Experiments have reported various other contrast enhancements: a factor of ${\sim}1.3$ in Ref. [46] and a factor of 1.5-1.6 in Ref. [49], though we cannot confirm if the microwave field amplitude was optimized in every case. The exact solution for arbitrary $T_{2}^{\ast}$, $\Omega_{R}$, $I_{0}$, and $T_{1}$ would require solving density-matrix quantum master equations [84, 85], which could make for an interesting future study.
@@ Appendix IV.4 Parallel ODMR fitting
The summand in Eq. (AVI-4) is a convolution of the $x$-component of the SPION dipole field, $B_{\rm dip,x}(x,y)$, with the microscope optical point-spread function, ${\rm PSF}(x,y)$. This expression is averaged over 5 different $z$ depths, ($z_{i}=-\{60,90,120,150,180\}~{}{\rm nm}$) selected to represent the relevant NV layer distribution, as described in Appendix II. The $60~{}{\rm nm}$ standoff is chosen to account for the SPION radius (${\sim}15~{}{\rm nm}$), a thin hydrocarbon layer between SPION and diamond (${\sim}10~{}{\rm nm}$), a (${\sim}15~{}{\rm nm}$) depleted layer near the diamond surface where implanted nitrogen is not efficiently converted to negatively-charged NV centers [36], and an additional (${\sim}20~{}{\rm nm}$) “dead layer” where the NV ODMR signal is not expected to contribute. The “dead layer” comes from the expectation that NV centers closer than ${\sim}60~{}{\rm nm}$ to a SPION center exhibit large enough ODMR line shifts that their contribution to the voxel-averaged ODMR lineshape is inadequately captured by a least-squares Lorentzian fit. Thus, while the NV layer is predicted to have a thickness of ${\sim}130~{}{\rm nm}$ (Appendix II), we sample the distribution in five planes extending over ${\sim}120~{}{\rm nm}$. This difference in NV distribution length has a relatively minor impact on the simulated magnetic field profiles, whereas other sources of uncertainty (for example, deviations from the expected SRIM profile) could play a larger role.
@@
Figure A13: Easy axis angle. An approximate value of $\cos{\theta}$ ) is given by $m_{x}/|\vec{m}|$, where $\theta$ is the angle of the easy axis with respect to the applied field direction and $\vec{m}$ is extracted from fits to SPIONs at the lowest practical applied field ($B_{0}=1.5~{}{\rm mT}$). A scatter plot of $m_{x}/|\vec{m}|$ vs. $B_{c}$ is shown for ${\sim}95$ SPIONs. As the standard expression for characteristic field is $B_{c}=k_{B}T/(\mu\cos{\theta})$ ), the approximately inverse relationship observed in this plot is expected.
@@
Appendix VIII.1 Bulk sample preparation and measurement
A Quantum Design MPMS-3 SQUID magnetometer was used to characterize SPIONs of the same batch as those imaged by diamond magnetic microscopy. Samples were prepared by suspending SPIONs in hexane, sonicating for approximately $10$ minutes to discourage particle aggregation, and depositing the suspension on a cotton swab. The cotton swab was fixed in a plastic straw sample holder for measurement inside the MPMS. An alternative sample preparation we used for bulk SPION characterization involved freeze-drying a SPION suspension in mannitol [87].
@@ Appendix VIII.2 Field-dependent magnetization curves
The mean of the single-SPION $m_{\rm sat}$ distribution [$\overline{m}_{\rm sat}=0.44~{}{\rm A{\cdot}nm^{2}}$, see Fig. 3(c)] can also be compared with literature values. Wu et al. [54] reported a saturation magnetic moment of $\overline{m}_{\rm sat}=2.5~{}{\rm A{\cdot}nm^{2}}$ per nanoparticle for 30-nm-diameter Ocean Nanotech SHA-30 SPIONs, and $\overline{m}_{\rm sat}=1.6~{}{\rm A{\cdot}nm^{2}}$ per nanoparticle for 25-nm-diameter Ocean Nanotech SHA-25 SPIONs. Ahrentorp et al. [53] reported $\overline{m}_{\rm sat}=2\mbox{-}3.4~{}{\rm A{\cdot}nm^{2}}$ for 25-nm SPIONs suspended in liquid. Mosavian et al. [36] used a super-resolution diamond magnetic microscopy technique to record $m_{x}=0.86~{}{\rm A{\cdot}nm^{2}}$ for a single Ocean Nanotech SOR30 SPION at an applied field of $30~{}{\rm mT}\,\hat{x}$. Our diamond magnetic microscopy measurements on 30-nm SPIONs yield $\overline{m}_{\rm sat}=0.44~{}{\rm A{\cdot}nm^{2}}$, which is somewhat smaller than these values.
There are several possible reasons why our method might record spuriously lower measured magnetic moments. Diamond magnetic microscopy can be affected by pixel cross-talk in which signals from distant NV centers contribute to the ODMR spectrum of a camera pixel near a SPION [36, 90]. Additionally, it is possible that the distribution of NV depths in the diamond is broader and deeper than estimated by SRIM; for example, due to ion channeling effects. This could be prevented by using delta-doped NV layers [69, 70, 71].
@@ Appendix VIII.3 AC susceptibility
The bulk AC susceptibility data is approximately log-normally distributed. The peak of the function fit to the 400 K data in A15(a) is $f_{\rm max}\approx 80~{}{\rm Hz}$, which corresponds to a modal value of $\tau_{N}\approx 2~{}{\rm ms}$. Taking the holding-field dependence $\tau_{N}=\tau_{N,0}\exp(-B_{\rm hold}/B_{N})$, as observed in Fig. 5(b) with $B_{N}\approx 0.7~{}{\rm mT}$, this translates to a modal value $\tau_{N}\approx 0.1~{}{\rm ms}$ at $B_{\rm hold}=2~{}{\rm mT}$. This peak value of $\tau_{N}$, together with a $\sim 6$-decade FWHM, give a log-normal distribution that approximately matches the experimental histogram data in Fig. 5(b) ($N=83$ for $\tau_{N}<0.06~{}{\rm s}$, $N=11$ for $0.06~{}{\rm s}<\tau_{N}<10~{}{\rm s}$, and $N=3$ for $\tau_{N}>10~{}{\rm s}$).
@@
Figures A17-A21 shows plots of the fitted values for $m_{x}$, $m_{y}$, and $m_{z}$ versus $B_{0}$ for each of the 101 SPIONs studied in Fig. 3 of the main text. The $m_{x}(B_{0})$ curves are fit to a Langevin function, see Eq. (1) of the main text. There are three features (numbered 109, 117, and 140) that are designated as outliers due to having fit results $B_{c}>7~{}{\rm mT}$ and $m_{\rm sat}>1.2~{}{\rm A{\cdot}nm^{2}}$. These outliers exhibit little-to-no saturation at the highest applied fields, so their Langevin fits are ambiguous.
@@
Figures A22-A24 show plots of $\Delta m(t)$ for 93 of the SPIONs in the field of view.